# Monotone Sequences of Metric Spaces with Compact Limits

### v1

@@ 4.7. Intrinsic Flat Convergence Theorem
First note that (4.106) follows from (4.84) combined with the monotonicity and Lemma 4.7.
To prove weak convergence, assume on the contrary that there is a tuple, $(\pi_{0},...\pi_{m})$, on $(Z,d_{Z})$, and a $\delta>0$ with a subsequence, we also denote $T_{j}$, such that
(4.107) $|T_{j}(\pi_{0},...\pi_{m})-T_{\infty}(\pi_{0},...\pi_{m})|<\delta\quad\forall j\in\mathbb{N}.$
Note that this subsequence $T_{j}$ satisfies the hypotheses of Proposition 4.17, so it has a subsequence which converges weakly to $T$, contradicting (4.107). ∎
4.10. Proving SWIF Convergence
In this section we complete the proof of Theorem 4.16:
@@ 4.10. Proving SWIF Convergence
By Corollary 4.18 of Proposition 4.17, we have pushforwards, $T_{j}=\chi_{j\#}T$, converging weakly to $T_{\infty}=\chi_{\infty\#}T$ as integral currents in $Z$.
By the uniform upper bound on total mass, the compactness of $(Z,d_{Z})$ and Wenger’s Flat=Weak Convergence theorem in [Wen07] (applied to metric spaces as explained in [SW11]) there exist integral currents, $B_{j},A_{j}$, on $(Z,d_{Z})$ such that
(4.108) $\partial B_{j}+A_{j}=T_{j}-T_{\infty}\textrm{ and }{\mathbf{M}}_{d_{Z}}(B_{j})+{\mathbf{M}}_{d_{Z}}(A_{j})\to 0$
so that
(4.109) $d_{F}^{Z}(\chi_{j\#}T,\chi_{\infty\#}T)\to 0.$
Taking $M_{\infty}=\operatorname{set}_{d_{\infty}}(T)$ and applying the fact that $\chi_{j}$ and $\chi_{\infty}$ are distance preserving maps, we have
(4.110) $d_{SWIF}((X,d_{j},T),(M_{\infty},d_{\infty},T))\to 0.$
Since we have the convergence of masses in Corollary 4.18, we have $\mathcal{VF}$ convergence as well. ∎
Remark 4.19.
It is an exercise to circumvent applying Wenger’s Flat=Weak Convergence Theorem in the above proof, by explicitly constructing a sequence of integral currents, $A_{j}$ and $B_{j}$, on $(Z,d_{Z})$ using the special properties of $(Z,d_{Z})$ proven in Prop 3.9. Since we have shown $T$ is an integral current on $(X,d_{\infty})$, there are integers, $a_{i}$, and charts
(4.111) $\psi_{i}:(U_{i}\subset{\mathbb{R}}^{m},d_{{\mathbb{R}}^{m}})\to(X,d_{\infty})\textrm{ with }\sum_{i=1}^{\infty}|a_{i}|\,\mathcal{H}^{m}_{d_{\infty}}(\psi_{i}(U_{i}))<\infty$
such that
(4.112) $T=\sum_{i=1}^{\infty}a_{i}\,\psi_{i\#}[[U_{i}]],$
and the same holds for $\partial T$. Then it is natural to define the integral current, $B_{j}$, using the same integers $a_{i}$ and charts,
(4.113) $\tilde{\psi}_{i,j}:(U_{i}\times[0,h_{j}]\subset{\mathbb{R}}^{m+1},d_{{\mathbb{R}}^{m+1}})\to(Z,d_{Z})$
such that
(4.114) $\tilde{\psi}_{i,J}(v,t)=\zeta_{j}(\psi_{i}(v),t).$
We then similarly construct the integral current, $A_{j}$, using the charts for $\partial T$ (which are not necessarily the boundaries of charts of $T$). Next we would verify that both atlases of charts are Lipschitz and that $A_{j}$ and $B_{j}$ have finite masses satisfying
(4.115) ${\mathbf{M}}_{d_{Z}}(A_{j})\leq\,C\,h_{j}\,{\mathbf{M}}_{d_{\infty}}(T)\textrm{ and }{\mathbf{M}}_{d_{Z}}(B_{j})\leq\,C\,h_{j}\,{\mathbf{M}}_{d_{\infty}}(T).$
Most difficult is showing $A_{j}+\partial B_{j}=T_{j}-T_{\infty}$ by texting against arbitrary Lipschitz tuples on $(Z,d_{Z})$. It may be necessary to restrict to $W_{j}$ to get everything to work.
@@ 5. Open Problems
We know that without the compactness assumption in our Theorem 1.1 and Theorem 4.16, that we may not have Gromov-Hausdorff Convergence and a common compact metric space. However we know that by Wenger’s Compactness Theorem the diameter and mass bounds imply that there is a subsequence converging to a SWIF limit,
(5.1) $(X,d_{j_{k}},T)\stackrel{{\scriptstyle\textrm{SWIF}}}{{\longrightarrow}}(X^{\prime}_{\infty},d^{\prime}_{\infty},T^{\prime}_{\infty}).$
Is it possible that with monotonicity, $d_{j}\leq d_{j+1}$, one can show any of the following:
  • the sequence itself converges,
  • there is volume preserving intrinsic flat convergence,
  • the limit, $X^{\prime}_{\infty}\subset X$, with restricted metric, $d^{\prime}_{\infty}=d_{\infty}$, possibly with closure, $\bar{X}_{\infty}=X$, and maybe $T_{\infty}=T$.
@@
  • The integral current structure, $T$, is assumed to be an integral current on $(X,d_{\infty})$ as an hypothesis, and that perhaps also that it has a bound on its multiplicity.

### v2

@@ 4.10. Proving SWIF Convergence
Note that these hypotheses imply the hypotheses of Proposition 3.7; so there exists a metric space
(4.111) $(Z_{a,b}=X\times[0,h],d_{Z_{a,b}})\textrm{ where }h=\epsilon/2,$
with
(4.112) $d_{Z_{a,b}}(z_{1},z_{2})\leq d_{taxi}(z_{1},z_{2})=d_{b}(x_{1},x_{2})+|t_{1}-t_{2}|,$
and there are distance preserving maps:
[…]
We will estimate the SWIF distance by constructing integral currents, $A$ and $B$, on $Z_{a,b}=X\times[0,h]$ such that
(4.115) $\partial B+A=f_{a\#}T_{a}-f_{b\#}T_{b}$
and then bounding the masses of $A$ and $B$.
We have an integral current space, $(M,d_{b},T_{b})$. So the integral current $T_{b}\times[0,h]$ is defined on the isometric product space, $(M_{b}\times[0,h],d_{isom})$ defined in Proposition 3.7 of [PS17] by Portegies-Sormani. This proposition states that
(4.116) $\partial(T_{b}\times[0,h])=-\partial T_{b}\times[[(0,h)]]+T_{b}\times\partial[[(0,h)]],$
and
(4.117) $T_{b}\times\partial[[(0,h)]]=\psi_{h\#}T_{b}-\psi_{0\#}T_{b}$
where $\psi_{t}(x)=(x,t)\in X\times[0,h]$ and
(4.118) ${\mathbf{M}}_{d_{isom}}(T_{b}\times[0,h])=h\,{\mathbf{M}}(T_{b}),$
where $d_{isom}$ is the standard isometric product:
(4.119) $d_{isom}((x_{1},t_{1}),(x_{2},t_{2}))=\sqrt{(d_{b}(x_{1},x_{2}))^{2}+(t_{1}-t_{2})^{2}}.$
We have $M_{b}\times[0,h]\subset X\times[0,h]=Z_{a,b}$, however $d_{Z_{a,b}})$ is not the standard isometric product distance. So we rescale the isometric product, compare to the taxi product, and apply (4.112) to see that:
(4.120) $\sqrt{2}\,d_{isom}(z_{1},z_{2})\geq d_{taxi}(z_{1},z_{2})\geq d_{Z_{a,b}}(z_{1},z_{2}).$
Note that $T_{b}\times[0,h]$ is an integral current on the rescaled isometric product space, $(X\times[0,h],\sqrt{2}d_{isom})$, satisfying (4.116)-(4.117) except now the mass rescales:
(4.121) ${\mathbf{M}}_{\sqrt{2}d_{isom}}(T_{b}\times[0,h])=2^{\frac{m+1}{2}}\,h{\mathbf{M}}_{d_{b}}(T_{b}).$
We can apply the Lipschitz one identity map,
(4.122) $\iota:(X\times[0,h],\sqrt{2}d_{isom})\to(X\times[0,h],\sqrt{2}d_{Z_{a,b}}),$
to define the pushforward integral current on $Z_{a,b}$:
(4.123) $B=\iota_{\#}(T_{b}\times[[(0,h)]])$
whose mass satisfies
(4.124) ${\mathbf{M}}_{d_{Z_{a,b}}}(B)\leq 2^{\frac{m+1}{2}}\,h\,{\mathbf{M}}_{d_{b}}(T_{b}).$
Similarly, we define
(4.125) $A=\iota_{\#}(\partial T_{b}\times[[(0,h)]])$
whose mass satisfies
(4.126) ${\mathbf{M}}_{d_{Z_{a,b}}}(A)\leq 2^{\frac{m}{2}}\,h\,{\mathbf{M}}_{d_{b}}(\partial T_{b}).$
By the pushforward of (4.116) we have
(4.127) $\partial B=-A+\iota_{\#}(\,T_{b}\times\partial[[(0,h)]]\,)$
and by the pushforward of (4.117) we have
(4.128) $\iota_{\#}(\,T_{b}\times\partial[[(0,h)]]\,)=\iota_{\#}\psi_{h\#}T_{b}-\iota_{\#}\psi_{0\#}T_{b}.$
Thus we need only show that
(4.129) $f_{a\#}T_{a}=\iota_{\#}\psi_{h\#}T_{b}$
and
(4.130) $f_{b\#}T_{b}=\iota_{\#}\psi_{0\#}T_{b}$
as integral currents on $(Z_{a,b},d_{Z_{a,b}})$. The latter easily follows from,
(4.131) $f_{b}(x)=(x,0)=\iota(x,0)=\iota(\psi_{0}(x)).$
We also have
(4.132) $f_{a}(x)=(x,h)=\iota(x,h)=\iota(\psi_{h}(x)).$
which gives
(4.133) $f_{a\#}T_{a}=\iota_{\#}\psi_{h\#}T_{a}.$
So we need only prove
(4.134) $\iota_{\#}\psi_{h\#}T_{a}=\iota_{\#}\psi_{h\#}T_{b}$
as integral currents on $(Z_{a,b},d_{Z_{a,b}})$.
Consider any Lipschitz tuple $(\pi_{0},...,\pi_{m})$ on $(Z_{a,b},d_{Z_{a,b}})$. Recall that
(4.135) $\iota_{\#}\psi_{h\#}T_{a}(\pi_{0},...,\pi_{m})=T_{a}(\pi_{0}\circ\iota\circ\psi_{h},...,\pi_{m}\circ\iota\circ\psi_{h}).$
For each Lipschitz function $\pi_{i}:(Z_{a,b},d_{Z_{a,b}})\to{\mathbb{R}}$ we claim that
(4.136) $\pi_{i}\circ\iota\circ\psi_{h}:(X,d_{a})\to{\mathbb{R}}\textrm{ is Lipschitz}.$
This follows from the fact that $f_{a}$ is distance preserving and so
(4.137) $|\pi_{i}(\iota(\psi_{h}(x_{1})))-\pi_{i}(\iota(\psi_{h}(x_{2})))|$ $=$ $|\pi_{i}(x_{1},h)-\pi_{i}(x_{2},h)|$
(4.138) $\leq$ $K\,d_{Z_{a,b}}((x_{1},h),(x_{2},h))$
(4.139) $=$ $K\,d_{Z_{a,b}}(f_{a}(x_{1}),f_{a}(x_{2}))$
(4.140) $=$ $K\,d_{a}(x_{1},x_{2}).$
So $(\pi_{0}\circ\iota\circ\psi_{h},...,\pi_{m}\circ\iota\circ\psi_{h})$ is a Lipschitz tuple on $(X,d_{a})$. Combining this with the hypothesis that “$T_{a}=T_{b}$ as integral currents on $(X,d_{a})$” we have
(4.141) $T_{a}(\pi_{0}\circ\iota\circ\psi_{h},...,\pi_{m}\circ\iota\circ\psi_{h})=T_{b}(\pi_{0}\circ\iota\circ\psi_{h},...,\pi_{m}\circ\iota\circ\psi_{h}).$
Thus, by the definition of pushforward, we have (4.134). ∎
We now apply Proposition 3.6, Proposition 4.17, and Proposition 4.18 to prove Theorem 4.16:
@@
First note that the hypotheses of Theorem 4.16 imply that the hypotheses of Proposition 3.6 and Proposition 4.17. By Proposition 3.6, we have uniform convergence of compact metric spaces $(X,d_{j})$ to a compact limit $(X,d_{\infty})$:
(4.142) $d_{\infty}(x,y)-\epsilon_{j}\leq d_{j}(x,y)\leq d_{\infty}(x,y)\quad\forall x,y\in X$
where $\epsilon_{j}\to 0$ as $j\to\infty.$
By Proposition 4.17 we have an integral current space,
(4.143) $(M_{\infty},d_{\infty},T_{\infty})\textrm{ where }M_{\infty}=\operatorname{set}(T_{\infty})\subset X$
with $Cl(M_{\infty})=X$ and “$T_{\infty}=T$ as integral currents on $(X,d_{j})$” for all $j\in{\mathbb{N}}$.
Next we apply Proposition 4.18 with
(4.144) $(X,d_{a},T_{a})=(X,d_{j},T)\textrm{ and }(X,d_{b},T_{b})=(X,d_{\infty},T_{\infty})$
where $\operatorname{set}(T_{j})=X$ and $\operatorname{set}(T_{\infty})=M_{\infty}$ and
(4.145) $h=h_{j}=2\epsilon_{j}\to 0\textrm{ as }j\to\infty.$
Thus by Proposition 4.18, we have
(4.146) $d_{SWIF}((X,d_{j},T),(M_{\infty},d_{\infty},T_{\infty}))\leq 2^{\frac{m+1}{2}}h_{j}({\mathbf{M}}_{d_{\infty}}(T_{\infty})+{\mathbf{M}}_{d_{\infty}}(\partial T_{\infty})).$
Thus by (4.84) of Proposition 4.17 we have,
(4.147) $d_{SWIF}((X,d_{j},T),(M_{\infty},d_{\infty},T_{\infty}))\leq 2^{\frac{m+1}{2}}\,h_{j}\,(V_{0}+A_{0})\to 0$
and we may conclude SWIF convergence of the original sequence.
Finally we apply the fact that ${\mathbf{M}}_{d_{j}}(T)$ is monotone increasing by Lemma 4.7 and the fact that a subsequence converges to ${\mathbf{M}}_{d_{\infty}}(T_{\infty})$ by Proposition 4.17 to conclude that mass converges. Thus we have volume preserving intrinsic flat convergence of the original sequence. ∎
@@ 5. Open Problems
Let us consider a monotone increasing sequence of integral current spaces, $(X,d_{j},T)$, with $d_{j}\leq d_{j+1}$ with the following three hypotheses
(5.1) $\operatorname{diam}_{d_{j}}(X)\leq D_{0},\quad{\mathbf{M}}_{d_{j}}(T)\leq V_{0},\quad{\mathbf{M}}_{d_{j}}(\partial T)\leq A_{0}.$
By Lemma 2.5, we have a pointwise limit, $(X,d_{\infty})$. We know that without the compactness assumption in our Theorem 1.1 and Theorem 4.16, that we may not have Gromov-Hausdorff Convergence of the sequence. See Example 2.6.
However, we know that by Wenger’s Compactness Theorem of [Wen11] that any sequence of integral current spaces satisfying (5.1) has a subsequence converging to a SWIF limit,
(5.2) $(X,d_{j_{k}},T)\stackrel{{\scriptstyle\textrm{SWIF}}}{{\longrightarrow}}(X^{\prime}_{\infty},d^{\prime}_{\infty},T^{\prime}_{\infty}),$
where the SWIF limit might be the zero space.
Combining the monotonicity assumptions with (5.1), perhaps one can show one or any of the following:
  • the sequence itself SWIF converges without needing a subsequence possibly to a nonzero limit space,
  • there is $\mathcal{VF}$ convergence to the limit space,
  • there is a distance preserving map from the SWIF limit, $(X^{\prime}_{\infty},d^{\prime}_{\infty})$, into the pointwise limit, $(X,d_{\infty})$, with possibly some relationship between the currents, $T_{\infty}^{\prime}$ and $T$.
@@
  • The integral current structure, $T$, is assumed to be an integral current on the pointwise limit, $(X,d_{\infty})$.